OpenStudy (anonymous):

Solve ΔABC with A = 61°, b = 33, and c = 39.

OpenStudy (dumbcow):

use law of cosines to find side "a" \[a^{2} = 33^{2}+39^{2}-2(33)(39)\cos 61\]

OpenStudy (dumbcow):

then find angle B using Law of sines \[\frac{\sin B}{33}=\frac{\sin 61}{a}\] then C = 180-61-B

OpenStudy (anonymous):

can you give me anymore help

OpenStudy (dumbcow):

well ive laid out the steps...the rest is just computing

OpenStudy (anonymous):

okay